OpenStudy (anonymous):

11. A catapult launches a boulder with an upward velocity of 148 ft/s. The height of the boulder, h, in feet after t seconds is given by the function . How long does it take the boulder to reach its maximum height? What is the boulder’s maximum height? Round to the nearest hundredth, if necessary. (1 point) 9.25 s; 30 ft 4.63 s; 640.5 ft 4.63 s; 1056.75 ft 4.63 s; 372.25 ft

OpenStudy (cwrw238):

at maxim height velocity = 0 use the formula v = u + at v=0,u=148, a = -32 0 = 148 - 32t 32t = 148 t = 148/32 = 4.625 seconds

OpenStudy (cwrw238):

thats 4.63 to nearest hundredth maximum height can be found by using the formula s = (u + v)t /2 = (148 + 0)*4.63/ 2 = 342.62 ft

OpenStudy (cwrw238):

which is not one of your options but i'm pretty sure i'm correct